OpenStudy (anonymous):

Paul can type 60 words per minute and Jennifer can type 80 words per minute. How does Paul's typing speed compare to Jennifer's? A. Paul can type 3/4 as fast as Jennifer. B. Paul can type 11/3 times as fast as Jennifer. C. Paul can type 1/3 as fast as Jennifer. D. Paul can type 1/4 as fast as Jennifer.

OpenStudy (anonymous):

A) Paul can type 3/4 times as fast as Jennifer, is the correct option

OpenStudy (anonymous):

well my typing speed compared to jeniffer is nothing but 60:80 which can also be written as 60/80 or 6/8 or 3/4

OpenStudy (austinl):

\(\dfrac{60}{80}=\dfrac{x}{y}\) Simplify the beginning fraction.

OpenStudy (austinl):

\(\dfrac{60}{80}=\dfrac{6\cancel{0}}{8\cancel{0}}=\dfrac{6}{8}\rightarrow \dfrac{\frac{6}{2}}{\frac{8}{2}}=\dfrac{3}{4}\)

OpenStudy (anonymous):

Paul can type is 3/4 times as fast as Jennifer (60/80 = 6/8 = 3/4)

OpenStudy (anonymous):

Thanks everyone (:

OpenStudy (austinl):

You are very welcome.